# Space-time FEM-BEM couplings for parabolic transmission problems

### v1

@@ 3.1. First-order system least squares formulation for interior problem
Note that ${\boldsymbol{u}}\cdot{\boldsymbol{n}}={\boldsymbol{\sigma}}\cdot{\boldsymbol{n}}_{\boldsymbol{x}}$ for all ${\boldsymbol{u}}=(u,{\boldsymbol{\sigma}})\in U$. By [24, Lemma 2.2],
@@
and (2.1) implies that $\mathcal{G}$ is indeed well defined and bounded. According to [24, Theorem 2.8], the latter two operators, $\mathcal{G}_{D}$ and $\mathcal{\mathcal{G}}_{N}$, are even bounded linear isomorphisms. Hence, $\mathcal{G}|_{U_{D}}$ and $\mathcal{G}|_{U_{N}}$ are bounded linear isomorphisms as well.
@@
$b_{\alpha}({\boldsymbol{u}},{\boldsymbol{v}}):=\alpha\langle\mathcal{G}{\boldsymbol{u}}\,,\,\mathcal{G}{\boldsymbol{v}}\rangle_{L}=\alpha\langle(f,\boldsymbol{0},u_{0})\,,\,\mathcal{G}{\boldsymbol{v}}\rangle_{L}\quad\text{for all }{\boldsymbol{v}}=(v,{\boldsymbol{\tau}})\in U.$ (3.3)
Here, $\langle\cdot\,,\,\cdot\rangle_{L}$ denotes the scalar product on $L$.
@@ 3.2.1. Dirichlet trace of representation formula
$\begin{split}b_{\alpha}({\boldsymbol{u}},{\boldsymbol{v}})+c_{1}({\boldsymbol{u}},{\boldsymbol{v}})=\alpha\langle(f,\boldsymbol{0},u_{0})\,,\,\mathcal{G}{\boldsymbol{v}}\rangle_{L}+\langle-\mathcal{V}\phi_{N}+(\mathcal{K}-\tfrac{1}{2})u_{D}\,,\,{\boldsymbol{\tau}}\cdot{\boldsymbol{n}}_{\boldsymbol{x}}\rangle_{\Sigma}\\ \text{for all }{\boldsymbol{v}}=(v,{\boldsymbol{\tau}})\in U.\end{split}$ (3.5)
@@ 3.2.2. Neumann trace of representation formula
$\begin{split}b_{\alpha}({\boldsymbol{u}},{\boldsymbol{v}})+c_{2}({\boldsymbol{u}},{\boldsymbol{v}})=\alpha\langle(f,\boldsymbol{0},u_{0})\,,\,\mathcal{G}{\boldsymbol{v}}\rangle_{L}+\langle\mathcal{W}u_{D}+(\mathcal{N}+\tfrac{1}{2})\phi_{N}\,,\,v|_{\Sigma}\rangle_{\Sigma}\\ \text{for all }{\boldsymbol{v}}=(v,{\boldsymbol{\tau}})\in U.\end{split}$ (3.7)
@@ 3.2.3. Dirichlet and Neumann trace of representation formula
$\begin{split}b_{\alpha}({\boldsymbol{u}},{\boldsymbol{v}})+c_{3}({\boldsymbol{u}},{\boldsymbol{v}})&=\alpha\langle(f,\boldsymbol{0},u_{0})\,,\,\mathcal{G}{\boldsymbol{v}}\rangle_{L}+\langle-\mathcal{V}\phi_{N}+(\mathcal{K}-\tfrac{1}{2})u_{D}\,,\,{\boldsymbol{\tau}}\cdot{\boldsymbol{n}}_{\boldsymbol{x}}\rangle_{\Sigma}\\ &\quad+\langle\mathcal{W}u_{D}+(\mathcal{N}+\tfrac{1}{2})\phi_{N}\,,\,v|_{\Sigma}\rangle_{\Sigma}\quad\text{for all }{\boldsymbol{v}}=(v,{\boldsymbol{\tau}})\in U.\end{split}$ (3.8)
@@ 3.3. Coercivity of combined bilinear forms
Note that $({\boldsymbol{A}}\nabla_{\boldsymbol{x}}u)\cdot{\boldsymbol{n}}_{\boldsymbol{x}}=-{\boldsymbol{\sigma}}\cdot{\boldsymbol{n}}_{\boldsymbol{x}}$. Integration by parts similar to [5, Equations (2.35)&(2.52)] together with $\partial_{t}u-\operatorname{\mathrm{div}}_{\boldsymbol{x}}({\boldsymbol{A}}\nabla_{\boldsymbol{x}}u)=0$ and $u(0,\cdot)=0$ gives that
@@
Next, we want to apply a standard compactness argument, more precisely, the Peetre–Tartar lemma [13, Lemma A.20]. In the notation from there (with $S$ instead of $T$, being already the end time point), we set
@@
With Lemma 3.2 and Lemma 3.3, coercivity of $b_{\alpha}+c_{i}$, $i\in\{1,2,3\}$, follows from the following abstract theorem.
@@ 3.4. Equivalence of couplings to transmission problem
$\begin{split}b_{\alpha}({\boldsymbol{u}},{\boldsymbol{v}})+c_{4}({\boldsymbol{u}},{\boldsymbol{v}})=\alpha\langle(f,\boldsymbol{0},u_{0})\,,\,\mathcal{G}{\boldsymbol{v}}\rangle_{L}+\langle\phi_{N}+\mathcal{V}^{-1}(\tfrac{1}{2}-\mathcal{K})u_{D}\,,\,v|_{\Sigma}\rangle_{\Sigma}\\ \text{for all }{\boldsymbol{v}}=(v,{\boldsymbol{\tau}})\in U.\end{split}$ (3.11)
@@
Elementary analysis and the stronger radiation condition $w({\boldsymbol{x}})={\mathcal{O}}(|x|^{-d})$ and $\nabla_{\boldsymbol{x}}w({\boldsymbol{x}})\cdot{\boldsymbol{x}}={\mathcal{O}}(|x|^{-d})$ as $|x|\to\infty$ from Remark 2.3 show that $\langle\nabla_{\boldsymbol{x}}w\cdot{\boldsymbol{n}}_{R,{\boldsymbol{x}}}\,,\,w|_{\Sigma_{R}}\rangle_{\Sigma_{R}}\to 0$ as $R\to\infty$; see also the proof of [5, Theorem 3.11]. The monotone convergence theorem yields that
@@ 4. Numerical experiments
$\begin{split}b_{\alpha}({\boldsymbol{u}}_{h},{\boldsymbol{v}}_{h})+c_{1}({\boldsymbol{u}}_{h},{\boldsymbol{v}}_{h})=\alpha\langle(f,\boldsymbol{0},u_{0})\,,\,\mathcal{G}{\boldsymbol{v}}_{h}\rangle_{L}+\langle-\mathcal{V}\phi_{N}+(\mathcal{K}-\tfrac{1}{2})u_{D}\,,\,{\boldsymbol{\tau}}_{h}\cdot{\boldsymbol{n}}_{\boldsymbol{x}}\rangle_{\Sigma}\end{split}$
@@ 4.4. Nonsmooth solutions
$\|\mathcal{G}({\boldsymbol{u}}-{\boldsymbol{u}}_{h})\|_{L}^{2}=\|(f,0,u_{0})-\mathcal{G}{\boldsymbol{u}}_{h}\|_{L}^{2},$
@@ Acknowledgement
TF acknowledges funding by ANID Chile through FONDECYT project 1210391. GG acknowledges funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC-2047/1 – 390685813. MK acknowledges funding by ANID Chile through FONDECYT project 1210579.

### v2

@@ 3.1. First-order system least squares formulation for interior problem
Note that ${\boldsymbol{u}}\cdot{\boldsymbol{n}}={\boldsymbol{\sigma}}\cdot{\boldsymbol{n}}_{\boldsymbol{x}}$ for all ${\boldsymbol{u}}=(u,{\boldsymbol{\sigma}})\in U$. By [28, Lemma 2.2],
@@
and (2.1) implies that $\mathcal{G}$ is indeed well defined and bounded. According to [28, Theorem 2.8], the latter two operators, $\mathcal{G}_{D}$ and $\mathcal{\mathcal{G}}_{N}$, are even bounded linear isomorphisms. Hence, $\mathcal{G}|_{U_{D}}$ and $\mathcal{G}|_{U_{N}}$ are bounded linear isomorphisms as well.
@@
$b_{\alpha}({\boldsymbol{u}},{\boldsymbol{v}}):=\alpha\langle\mathcal{G}{\boldsymbol{u}}\,,\,\mathcal{G}{\boldsymbol{v}}\rangle_{L}=\alpha\langle(f,\boldsymbol{0},u_{0})\,,\,\mathcal{G}{\boldsymbol{v}}\rangle_{L}\quad\text{for all }{\boldsymbol{v}}=(v,{\boldsymbol{\tau}})\in U,$ (3.3)
Here, $\langle\cdot\,,\,\cdot\rangle_{L}$ denotes the scalar product on $L$. More explicitly, this reads as
$b_{\alpha}({\boldsymbol{u}},{\boldsymbol{v}})$ $=\alpha\big{[}\langle\operatorname{\mathrm{div}}{\boldsymbol{u}}\,,\,\operatorname{\mathrm{div}}{\boldsymbol{v}}\rangle_{L^{2}(Q)}+\langle{\boldsymbol{A}}\nabla_{\boldsymbol{x}}u+{\boldsymbol{\sigma}}\,,\,{\boldsymbol{A}}\nabla_{\boldsymbol{x}}v+{\boldsymbol{\tau}}\rangle_{L^{2}(Q)}+\langle u(0,\cdot)\,,\,v(0,\cdot)\rangle_{L^{2}(\Omega)}\big{]}$
$=\alpha\big{[}\langle f\,,\,\operatorname{\mathrm{div}}{\boldsymbol{v}}\rangle_{L^{2}(Q)}+\langle u_{0}\,,\,v(0,\cdot)\rangle_{L^{2}(\Omega)}\big{]},$
where $\langle\cdot\,,\,\cdot\rangle_{L^{2}(Q)}$ denotes the scalar product on $L^{2}(Q)$ and $\langle\cdot\,,\,\cdot\rangle_{L^{2}(\Omega)}$ the scalar product on $L^{2}(\Omega)$.
@@ 3.2.1. Dirichlet trace of representation formula
$b_{\alpha}({\boldsymbol{u}},{\boldsymbol{v}})+c_{1}({\boldsymbol{u}},{\boldsymbol{v}})=\alpha\big{[}\langle f\,,\,\operatorname{\mathrm{div}}{\boldsymbol{v}}\rangle_{L^{2}(Q)}+\langle u_{0}\,,\,v(0,\cdot)\rangle_{L^{2}(\Omega)}\big{]}+\langle-\mathcal{V}\phi_{N}+(\mathcal{K}-\tfrac{1}{2})u_{D}\,,\,{\boldsymbol{\tau}}\cdot{\boldsymbol{n}}_{\boldsymbol{x}}\rangle_{\Sigma}$
for all $\hskip 312.9803pt\text{for all }{\boldsymbol{v}}=(v,{\boldsymbol{\tau}})\in U.$ (3.5)
@@ 3.2.2. Neumann trace of representation formula
$b_{\alpha}({\boldsymbol{u}},{\boldsymbol{v}})+c_{2}({\boldsymbol{u}},{\boldsymbol{v}})=\alpha\big{[}\langle f\,,\,\operatorname{\mathrm{div}}{\boldsymbol{v}}\rangle_{L^{2}(Q)}+\langle u_{0}\,,\,v(0,\cdot)\rangle_{L^{2}(\Omega)}\big{]}+\langle\mathcal{W}u_{D}+(\mathcal{N}+\tfrac{1}{2})\phi_{N}\,,\,v|_{\Sigma}\rangle_{\Sigma}$
for all $\hskip 312.9803pt\text{for all }{\boldsymbol{v}}=(v,{\boldsymbol{\tau}})\in U.$ (3.7)
@@ 3.2.3. Dirichlet and Neumann trace of representation formula
$b_{\alpha}({\boldsymbol{u}},{\boldsymbol{v}})+c_{3}({\boldsymbol{u}},{\boldsymbol{v}})=\alpha\big{[}\langle f\,,\,\operatorname{\mathrm{div}}{\boldsymbol{v}}\rangle_{L^{2}(Q)}+\langle u_{0}\,,\,v(0,\cdot)\rangle_{L^{2}(\Omega)}\big{]}+\langle-\mathcal{V}\phi_{N}+(\mathcal{K}-\tfrac{1}{2})u_{D}\,,\,{\boldsymbol{\tau}}\cdot{\boldsymbol{n}}_{\boldsymbol{x}}\rangle_{\Sigma}$
$\hskip 142.26378pt+\langle\mathcal{W}u_{D}+(\mathcal{N}+\tfrac{1}{2})\phi_{N}\,,\,v|_{\Sigma}\rangle_{\Sigma}\quad\text{for all }{\boldsymbol{v}}=(v,{\boldsymbol{\tau}})\in U.$ (3.8)
@@ 3.3. Coercivity of combined bilinear forms
Note that $({\boldsymbol{A}}\nabla_{\boldsymbol{x}}u)\cdot{\boldsymbol{n}}_{\boldsymbol{x}}=-{\boldsymbol{\sigma}}\cdot{\boldsymbol{n}}_{\boldsymbol{x}}$. Integration by parts similar to [7, Equations (2.35)&(2.52)] together with $\partial_{t}u-\operatorname{\mathrm{div}}_{\boldsymbol{x}}({\boldsymbol{A}}\nabla_{\boldsymbol{x}}u)=0$ and $u(0,\cdot)=0$ gives that
@@
Next, we want to apply a standard compactness argument, more precisely, the Peetre–Tartar lemma [16, Lemma A.20]. In the notation from there (with $S$ instead of $T$, being already the end time point), we set
@@
With Lemma 3.2 and Lemma 3.3, coercivity of $b_{\alpha}+c_{i}$, $i\in\{1,2,3\}$, follows from the following abstract theorem, which is also implicitly used in [23].
@@ 3.4. Equivalence of couplings to transmission problem
$b_{\alpha}({\boldsymbol{u}},{\boldsymbol{v}})+c_{4}({\boldsymbol{u}},{\boldsymbol{v}})=\alpha\big{[}\langle f\,,\,\operatorname{\mathrm{div}}{\boldsymbol{v}}\rangle_{L^{2}(Q)}+\langle u_{0}\,,\,v(0,\cdot)\rangle_{L^{2}(\Omega)}\big{]}+\langle\phi_{N}+\mathcal{V}^{-1}(\tfrac{1}{2}-\mathcal{K})u_{D}\,,\,v|_{\Sigma}\rangle_{\Sigma}$
for all $\hskip 312.9803pt\text{for all }{\boldsymbol{v}}=(v,{\boldsymbol{\tau}})\in U.$ (3.11)
@@
Elementary analysis and the stronger radiation condition $w({\boldsymbol{x}})={\mathcal{O}}(|x|^{-d})$ and $\nabla_{\boldsymbol{x}}w({\boldsymbol{x}})\cdot{\boldsymbol{x}}={\mathcal{O}}(|x|^{-d})$ as $|x|\to\infty$ from Remark 2.3 show that $\langle\nabla_{\boldsymbol{x}}w\cdot{\boldsymbol{n}}_{R,{\boldsymbol{x}}}\,,\,w|_{\Sigma_{R}}\rangle_{\Sigma_{R}}\to 0$ as $R\to\infty$; see also the proof of [7, Theorem 3.11]. The monotone convergence theorem yields that
@@ 4. Numerical experiments
$\begin{split}b_{\alpha}({\boldsymbol{u}}_{h},{\boldsymbol{v}}_{h})+c_{1}({\boldsymbol{u}}_{h},{\boldsymbol{v}}_{h})=\alpha\big{[}\langle f\,,\,\operatorname{\mathrm{div}}{\boldsymbol{v}}\rangle_{L^{2}(Q)}+\langle u_{0}\,,\,v(0,\cdot)\rangle_{L^{2}(\Omega)}\big{]}\\ +\langle-\mathcal{V}\phi_{N}+(\mathcal{K}-\tfrac{1}{2})u_{D}\,,\,{\boldsymbol{\tau}}_{h}\cdot{\boldsymbol{n}}_{\boldsymbol{x}}\rangle_{\Sigma}\end{split}$
@@ 4.4. Nonsmooth solutions
$\|\mathcal{G}({\boldsymbol{u}}-{\boldsymbol{u}}_{h})\|_{L}^{2}$ $=\|(f,0,u_{0})-\mathcal{G}{\boldsymbol{u}}_{h}\|_{L}^{2}$
$=\|f-\operatorname{\mathrm{div}}{\boldsymbol{u}}_{h}\|_{L^{2}(Q)}^{2}+\|{\boldsymbol{A}}\nabla_{\boldsymbol{x}}u_{h}+{\boldsymbol{\sigma}}_{h}\|_{L^{2}(Q)}^{2}+\|u_{0}-u_{h}(0,\cdot)\|_{L^{2}(\Omega)}^{2},$
@@ Acknowledgement
TF acknowledges funding by ANID Chile through FONDECYT project 1210391 & project 1250070. GG acknowledges funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC-2047/1 – 390685813. MK acknowledges funding by ANID Chile through FONDECYT project 1210579 & project 1250604.